OpenStudy (anonymous):

What is the third-degree polynomial function such that f(0) = -12 and whose zeros are 1, 2, and 3

OpenStudy (anonymous):

answers: f(x) = 4x3 - 24x2 + 22x - 24 f(x) = 2x3 - 12x2 + 22x - 12 f(x) = 4x3 + 24x2 - 44x - 24 f(x) = 2x3 + 12x2 - 22x - 12

OpenStudy (precal):

y=a(x-1)(x-2)(x-3) I suppose you have to sub x=0 and y=-12 to get your a value and then multiply it out